The average value of all the pennies, nickels, dimes and quarters in Bob's pocket is 20 cents. If he had one more quarter, the average would be 21 cents. How many dimes does he have?
Many thanks!
: The average value of all the pennies, nickels, dimes and quarters in Bob's pocket is 20 cents. If he had one more quarter, the average would be 21 cents. How many dimes does he have?
Many thanks!

Answer by scott8148(2758)

working in cents

let n=number of coins

"The average value of all the pennies, nickels, dimes and quarters in Bob's pocket is 20 cents" __ 20n

"If he had one more quarter, the average would be 21 cents" __ 20n+25=21(n+1)

20n+25=21n+21 __ 4=n

4 coins worth 80 cents means 3 quarters and a nickel
__ no pennies or dimes